# Fisher-Rao Gradient Flow: Geodesic Convexity and Functional Inequalities

### v1

@@ 4.2. Sufficient and necessary conditions for gradient dominance
(4.6) $\begin{split}\bigl{(}f^{\prime}(y)-f^{\prime}(x)\bigr{)}^{2}\geq\alpha(\frac{1}{x}-\frac{1}{y})\bigl{(}\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\bigr{)}\end{split}$
@@
Given a convex function $f\in C^{2}(0,+\infty)$ with $f(1)=0$ and $f^{\prime}(1)=0$, then Condition 4.2 is sufficient and necessary for gradient dominance (4.1) to hold.
Sketch of Proof.
The proof breaks into the following three steps. In the first step, we use measure theoretical tools to reduce the gradient dominance condition to an inequality on the discrete probability simplex, which is simpler to deal with. In the second step, we use novel linear programming techniques to reduce the global nonlocal term in the inequality to a three-point interaction term. Finally, we show the two-point version of inequality, which is the sufficient and necessary condition, implies the three-point ones, thus completing the proof. We present a sketch of main techniques here and refer to the Appendix for detailed proofs of the arguments.
@@
There exists a sequence of bounded domains $\Omega_{n}\subset\mathbb{R}^{d}$ for $n\geq 1$ such that
(4.7) $\Omega_{n}\subset\Omega_{n+1},\quad\bigcup_{n\geq 1}\Omega_{n}=\mathbb{R}^{d},\quad\text{and}\quad\int_{\Omega_{n}}\rho(\theta){\rm d}\theta=\int_{\Omega_{n}}\rho^{*}(\theta){\rm d}\theta.$
@@
(4.8) $\rho_{n}(\theta)=\frac{1}{\int_{\Omega_{n}}\rho(\theta^{\prime}){\rm d}\theta^{\prime}}\rho(\theta)\mathbbm{1}_{\Omega_{n}}(\theta)\quad\text{and}\quad\rho^{*}_{n}(\theta)=\frac{1}{\int_{\Omega_{n}}\rho^{*}(\theta^{\prime}){\rm d}\theta^{\prime}}\rho^{*}(\theta)\mathbbm{1}_{\Omega_{n}}(\theta).$
@@
(4.9) $\int\rho_{n}\bigl{(}f^{\prime}\bigl{(}\frac{\rho_{n}}{\rho^{*}_{n}}\bigr{)}-c\bigr{)}^{2}\mathrm{d}\theta\geq\alpha_{f}D_{f}[\rho_{n}\|\rho^{*}_{n}],$
for any constant $c\in\mathbb{R}$. The rationale behind this claim is that, once the above hold, we can take $n\to\infty$. Here we have
(4.10) $\lim_{n\to\infty}\int\rho_{n}\big{(}f^{\prime}\bigl{(}\frac{\rho_{n}}{\rho^{*}_{n}}\bigr{)}-c\big{)}^{2}\mathrm{d}\theta=\frac{\lim_{n\to\infty}\int_{\Omega_{n}}\rho\big{(}f^{\prime}\bigl{(}\frac{\rho}{\rho^{*}}\bigr{)}-c\big{)}^{2}\mathrm{d}\theta}{\lim_{n\to\infty}\int_{\Omega_{n}}\rho(\theta){\rm d}\theta}=\int\rho\bigl{(}f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}-c\bigr{)}^{2},$
@@
(4.11) $\int\rho\bigl{(}f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}-c\bigr{)}^{2}\geq\alpha D_{f}[\rho\|\rho^{*}].$
By optimizing the constant $c$, we arrive at the desired gradient dominance condition.
Therefore, it suffices to prove the claim in (4.9), which is the same inequality as (4.1) but applied to densities in bounded domains. Suppose $\Omega_{n}=\cup_{k=1}^{K}\omega_{k}$ is partitioned to $K$ disjoint subdomains. We can approximate $\rho_{n}$ and $\rho^{*}$ by
$\sum_{k=1}^{K}\frac{\int_{\omega_{k}}\rho_{n}{\rm d}\theta}{\int_{\omega_{k}}{\rm d}\theta}\mathbbm{1}_{\omega_{k}}\quad\text{and}\quad\sum_{k=1}^{K}\frac{\int_{\omega_{k}}\rho_{n}^{*}{\rm d}\theta}{\int_{\omega_{k}}{\rm d}\theta}\mathbbm{1}_{\omega_{k}}.$
If we can prove the inequality for the above approximate densities, we can then take finer and finer partitions and as $K\to\infty$, we will get the corresponding inequality for $\rho_{n}$ and $\rho^{*}$. This is true because we work in a bounded domain $\Omega_{n}$, for which we can apply uniform convergence to establish the limit; note that $\rho$ and $\rho^{*}$ are assumed smooth.
@@
(4.12) $\sum_{i=1}^{K}\rho_{i}\bigl{(}f^{\prime}(x_{i})-\sum_{j=1}^{K}\rho_{j}f^{\prime}(x_{j})\bigr{)}^{2}\geq\alpha_{f}\sum_{i}\frac{\rho_{i}}{x_{i}}f(x_{i}),$
@@
(4.13) $\sum_{i=1}^{K}\rho_{i}=1,\quad\sum_{i=1}^{K}\frac{\rho_{i}}{x_{i}}=1,\quad\rho_{i}>0,\quad x_{i}>0.$
@@
(4.14) $\sum_{i=1}^{K}\rho_{i}\bigl{(}f^{\prime}(x_{i})-\sum_{j=1}^{K}\rho_{j}f^{\prime}(x_{j})\bigr{)}^{2}=\sum_{i=1}^{K}\rho_{i}(f^{\prime}(x_{i}))^{2}-(\sum_{j=1}^{K}\rho_{j}f^{\prime}(x_{j}))^{2}$
$=$ $\frac{1}{2}\left(\sum_{i=1}^{K}\rho_{i}(f^{\prime}(x_{i}))^{2}+\sum_{i=j}^{K}\rho_{j}(f^{\prime}(x_{j}))^{2}-2(\sum_{i=1}^{K}\rho_{j}f^{\prime}(x_{i}))(\sum_{j=1}^{K}\rho_{j}f^{\prime}(x_{j}))\right)$
@@
(4.15) $\begin{split}\sum_{i=1}^{K}\frac{\rho_{i}}{x_{i}}f(x_{i})&=\sum_{i=1}^{K}\frac{\rho_{i}}{x_{i}(1-x_{i})}f(x_{i})-\sum_{i}\frac{\rho_{i}}{1-x_{i}}f(x_{i})\\ &=\sum_{1\leq i,j\leq K}\frac{\rho_{i}\rho_{j}}{x_{i}(1-x_{i})}f(x_{i})-\sum_{ij}\frac{\rho_{i}\rho_{j}}{x_{j}(1-x_{i})}f(x_{i})\\ &=\frac{1}{2}\sum_{1\leq i,j\leq K}\rho_{i}\rho_{j}(\frac{1}{x_{j}}-\frac{1}{x_{i}})\bigl{(}\frac{f(x_{j})}{1-x_{j}}-\frac{f(x_{i})}{1-x_{i}}\bigr{)},\end{split}$
@@
where to get the second line we used (4.13). We remark that since $f\in C^{2}(0,+\infty)$ and $f(1)=0$, $f(x)/(1-x)$ is well defined and there is no singularity at $x=1$. Therefore, the equivalent inequality is
(4.16) $\sum_{1\leq i,j\leq K}\rho_{i}\rho_{j}\Bigl{(}f^{\prime}(x_{i})-f^{\prime}(x_{j})\Bigr{)}^{2}\geq\alpha_{f}\sum_{1\leq i,j\leq K}\rho_{i}\rho_{j}(\frac{1}{x_{i}}-\frac{1}{x_{j}})\Bigl{(}\frac{f(x_{i})}{x_{i}-1}-\frac{f(x_{j})}{x_{j}-1}\Bigr{)}.$
@@
When $K=2$, denote $x=x_{1}$ and $y=x_{2}$ with $0<x<1<y$. We have $\rho_{1}=\frac{x(y-1)}{y-x}$ and $\rho_{2}=\frac{y(1-x)}{y-x}$, (4.16) leads to
@@
(4.17) $\begin{split}\exists\ \alpha>0\textrm{ s.t. }&\bigl{(}f^{\prime}(y)-f^{\prime}(x)\bigr{)}^{2}\geq\alpha(\frac{1}{x}-\frac{1}{y})\bigl{(}\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\bigr{)},\ \forall\ 0<x<1<y\end{split}$
@@
with $\alpha=\alpha_{f}$. This is the same as Condition 4.2. It is a necessary condition for the discrete inequality, and is also necessary for the continuous inequality by smoothing the indicator functions. It remains to show that Condition 4.2 is also sufficient, i.e., the case $K=2$ holds implies that the inequality holds for any $K\geq 3$.
@@
We first show that cases $K\geq 3$ can be reduced to $K=3$. To do so, we use the original formulation (4.12). We fix $x_{i}>0$, $\alpha_{f}>0$, and $M=\sum_{i=1}^{K}\rho_{i}f^{\prime}(x_{i})$, and consider the following linear programming problem with respect to variables $\{\rho_{i}\}_{i=1}^{K}$
$\min$ $\sum_{i=1}^{K}\rho_{i}\Bigl{(}f^{\prime}(x_{i})^{2}-\alpha_{f}\frac{f(x_{i})}{x_{i}}\Bigr{)}-M^{2}$
s.t. $\sum_{i=1}^{K}\rho_{i}=1,\ \sum_{i=1}^{K}\frac{\rho_{i}}{x_{i}}=1,\ \sum_{i=1}^{K}\rho_{i}f^{\prime}(x_{i})=M,\ \rho_{i}\geq 0.$
@@
The linear programming problem has a feasible solution and the constraint set is bounded. We claim that there is a minimizer that has at most $3$ nonzero $\rho_{i}$. Otherwise, without loss of generality, we assume for the minimizer $\rho$, it holds $\rho_{j}>0$ for $1\leq j\leq 4$. We can adjust them in the direction $\vec{d}=[d_{1},\,d_{2},\,d_{3},\,d_{4}]$ satisfying $\sum_{j=1}^{K}d_{j}=0$, $\sum_{j=1}^{K}\frac{d_{j}}{x_{j}}=0$, and $\sum_{j=1}^{K}d_{j}f^{\prime}(x_{j})=0$ to reduce the objective function or increase the number of nonzero $\rho_{i}$. This leads to a contradition so the claim is true. This means that we only need to prove the inequality for the case $K=3$. Using the equivalent formulation (4.16), assuming (4.6) holds, we need to show
(4.18) $\sum_{1\leq i,j\leq 3}\rho_{i}\rho_{j}\Bigl{(}f^{\prime}(x_{i})-f^{\prime}(x_{j})\Bigr{)}^{2}\geq\alpha_{f}\sum_{1\leq i,j\leq 3}\rho_{i}\rho_{j}(\frac{1}{x_{i}}-\frac{1}{x_{j}})\Bigl{(}\frac{f(x_{i})}{x_{i}-1}-\frac{f(x_{j})}{x_{j}-1}\Bigr{)}.$
@@
The last step is to use the inequality in the case of $K=2$ to prove (4.18). We have the following two technical lemmas whose proofs are in Appendix A.
@@
Assume (4.6) holds. There exists $10<\delta<$ 1, and $0<\alpha_{\delta}\leq\alpha$, such that
(4.19) $\bigl{(}f^{\prime}(y)-f^{\prime}(x)\bigr{)}^{2}\geq\alpha_{\delta}(\frac{1}{x}-\frac{1}{y})\bigl{(}\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\bigr{)},\ \forall\ x\in(1-\delta,1+\delta)\ \textrm{and}\ y>0.$
@@
Given a convex function $f\in C^{2}(0,+\infty)$ with $f(1)=0$ and $f^{\prime}(1)=0$ and under the constraint (4.13), conditions (4.6) and (4.19) implies (4.18).
With the above lemmas, the sufficient condition part is proved. ∎
4.3. An easier-to-verify sufficient condition
We provide a simpler and easier-to-verify sufficient condition on $f$ to ensure the gradient dominance 4.1 holds.
Theorem 4.6.
We assume $f\in C^{2}(0,+\infty)$ is a convex function with $f(1)=0$ and there exists $\alpha_{s}>0$ such that $x^{2}f^{\prime\prime}(x)>\alpha_{s}$ for $0<x\leq 1$. Then Condition 4.2 holds. Note that the assumptions are satisfied with $f^{\prime\prime}(x)=x^{p}$ and $p\leq-2$.
Proof.
The assumption $x^{2}f^{\prime\prime}(x)>\alpha_{s}$ for $0<x\leq 1$ implies that $f^{\prime\prime}(1)>\alpha_{s}$. Combining this with the condition $f\in C^{2}$ leads to the existence of $0<\delta<\frac{1}{2}$, such that
(4.20) $\frac{f^{\prime\prime}(1)}{2}\leq f^{\prime\prime}(x)\quad\text{for}\quad 1-2\delta\leq x\leq 1+2\delta.$

### v2

@@ 4.2 Sufficient and necessary conditions for gradient dominance
(59) $\begin{split}\Bigl(f^{\prime}(y)-f^{\prime}(x)\Bigr)^{2}\geq\alpha(\frac{1}{x}-\frac{1}{y})\Bigl(\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\Bigr)\end{split}$
@@
Given a convex function $f\in C^{2}(0,+\infty)$ with $f(1)=0$ and $f^{\prime}(1)=0$, then 4.2 is sufficient and necessary for gradient dominance (4.1) to hold.
Proof 4.4 (Sketch of Proof).
The proof breaks into the following three steps. In the first step, we use measure theoretical tools to reduce the gradient dominance condition to an inequality on the discrete probability simplex, which is simpler to deal with. In the second step, we use novel linear programming techniques to reduce the global nonlocal term in the inequality to a three-point interaction term. Finally, we show the two-point version of inequality, which is the sufficient and necessary condition, implies the three-point ones, thus completing the proof. We present a sketch of the main techniques here and refer to the Appendix for detailed proofs of the arguments.
@@
Given smooth densities $\rho$ and $\rho^{*}$ defined on $\mathbb{R}^{d}$, there exists a sequence of bounded domains $\Omega_{n}\subset\mathbb{R}^{d}$ for $n\geq 1$ such that
(60) $\Omega_{n}\subset\Omega_{n+1},\quad\bigcup_{n\geq 1}\Omega_{n}=\mathbb{R}^{d},\quad\text{and}\quad\int_{\Omega_{n}}\rho(\theta){\rm d}\theta=\int_{\Omega_{n}}\rho^{*}(\theta){\rm d}\theta.$
@@
(61) $\rho_{n}(\theta)=\frac{1}{\int_{\Omega_{n}}\rho(\theta^{\prime}){\rm d}\theta^{\prime}}\rho(\theta)\mathbbm{1}_{\Omega_{n}}(\theta)\quad\text{and}\quad\rho^{*}_{n}(\theta)=\frac{1}{\int_{\Omega_{n}}\rho^{*}(\theta^{\prime}){\rm d}\theta^{\prime}}\rho^{*}(\theta)\mathbbm{1}_{\Omega_{n}}(\theta).$
@@
(62) $\int\rho_{n}\Bigl(f^{\prime}\bigl(\frac{\rho_{n}}{\rho^{*}_{n}}\bigr)-c\Bigr)^{2}\mathrm{d}\theta\geq\alpha_{f}D_{f}[\rho_{n}\|\rho^{*}_{n}],$
for any constant $c\in\mathbb{R}$. The rationale behind this claim is that, once the above holds, we can take $n\to\infty$. Here we have
(63) $\lim_{n\to\infty}\int\rho_{n}\Big(f^{\prime}\bigl(\frac{\rho_{n}}{\rho^{*}_{n}}\bigr)-c\Big)^{2}\mathrm{d}\theta=\frac{\lim_{n\to\infty}\int_{\Omega_{n}}\rho\Big(f^{\prime}\bigl(\frac{\rho}{\rho^{*}}\bigr)-c\Big)^{2}\mathrm{d}\theta}{\lim_{n\to\infty}\int_{\Omega_{n}}\rho(\theta){\rm d}\theta}=\int\rho\Bigl(f^{\prime}\bigl(\frac{\rho}{{\rho^{*}}}\bigr)-c\Bigr)^{2},$
@@
(64) $\int\rho\Bigl(f^{\prime}\bigl(\frac{\rho}{{\rho^{*}}}\bigr)-c\Bigr)^{2}\geq\alpha_{f}D_{f}[\rho\|\rho^{*}].$
By setting $c=\mathbb{E}_{\rho}[f^{\prime}(\frac{\rho}{\rho^{*}})]$ to achieve the minimum on the left hand side, we arrive at the desired gradient dominance condition.
Therefore, it suffices to prove the claim in (62), which is the same inequality as (55) but applied to densities in bounded domains. Suppose $\Omega_{n}=\cup_{k=1}^{K}\omega_{k}$ is partitioned to $K$ disjoint subdomains. We can approximate $\rho_{n}$ and $\rho_{n}^{*}$ by $\sum_{k=1}^{K}\frac{\int_{\omega_{k}}\rho_{n}{\rm d}\theta}{\int_{\omega_{k}}{\rm d}\theta}\mathbbm{1}_{\omega_{k}}$ and $\quad\sum_{k=1}^{K}\frac{\int_{\omega_{k}}\rho_{n}^{*}{\rm d}\theta}{\int_{\omega_{k}}{\rm d}\theta}\mathbbm{1}_{\omega_{k}}$. If we can prove the inequality for the above approximate densities, we can then take finer and finer partitions and as $K\to\infty$, we will get the corresponding inequality for $\rho_{n}$ and $\rho^{*}$. This is true because we work in a bounded domain $\Omega_{n}$, for which we can apply uniform convergence to establish the limit; note that $\rho$ and $\rho^{*}$ are assumed smooth.
@@
(65) $\sum_{i=1}^{K}\rho_{i}\Bigl(f^{\prime}(x_{i})-\sum_{j=1}^{K}\rho_{j}f^{\prime}(x_{j})\Bigr)^{2}\geq\alpha_{f}\sum_{i}\frac{\rho_{i}}{x_{i}}f(x_{i}),$
@@
(66) $\sum_{i=1}^{K}\rho_{i}=1,\quad\sum_{i=1}^{K}\frac{\rho_{i}}{x_{i}}=1,\quad\rho_{i}>0,\quad x_{i}>0.$
@@
(67) $\sum_{i=1}^{K}\rho_{i}\bigl(f^{\prime}(x_{i})-\sum_{j=1}^{K}\rho_{j}f^{\prime}(x_{j})\bigr)^{2}=\sum_{i=1}^{K}\rho_{i}\bigl(f^{\prime}(x_{i})\bigr)^{2}-\bigl(\sum_{j=1}^{K}\rho_{j}f^{\prime}(x_{j})\bigr)^{2}$
$=$ $\frac{1}{2}\left(\sum_{i=1}^{K}\rho_{i}\bigl(f^{\prime}(x_{i})\bigr)^{2}+\sum_{j=1}^{K}\rho_{j}\bigl(f^{\prime}(x_{j})\bigr)^{2}-2\bigl(\sum_{i=1}^{K}\rho_{j}f^{\prime}(x_{i})\bigr)\bigl(\sum_{j=1}^{K}\rho_{j}f^{\prime}(x_{j})\bigr)\right)$
@@
(68) $\begin{split}\sum_{i=1}^{K}\frac{\rho_{i}}{x_{i}}f(x_{i})&=\sum_{i=1}^{K}\frac{\rho_{i}}{x_{i}(1-x_{i})}f(x_{i})-\sum_{i=1}^{K}\frac{\rho_{i}}{1-x_{i}}f(x_{i})\\ &=\sum_{1\leq i,j\leq K}\rho_{j}\frac{\rho_{i}}{x_{i}(1-x_{i})}f(x_{i})-\sum_{1\leq i,j\leq K}\frac{\rho_{j}}{x_{j}}\frac{\rho_{i}}{1-x_{i}}f(x_{i})\\ &=\sum_{1\leq i,j\leq K}\rho_{i}\rho_{j}(\frac{1}{x_{i}}-\frac{1}{x_{j}})\frac{f(x_{i})}{1-x_{i}}\\ &=\frac{1}{2}\sum_{1\leq i,j\leq K}\rho_{j}\rho_{i}(\frac{1}{x_{j}}-\frac{1}{x_{i}})\Bigl(\frac{f(x_{j})}{1-x_{j}}-\frac{f(x_{i})}{1-x_{i}}\Bigr),\end{split}$
@@
where to get the second line we used (66). We remark that since $f\in C^{2}(0,+\infty)$ and $f(1)=0$, f(x)/(1x)f(x)/(1-x) is well defined and there is no singularity at $x=1$ (we define the value here through the limit $x\to 1$). Therefore, the equivalent inequality is
(69) $\sum_{1\leq i,j\leq K}\rho_{i}\rho_{j}\Bigl(f^{\prime}(x_{i})-f^{\prime}(x_{j})\Bigr)^{2}\geq\alpha_{f}\sum_{1\leq i,j\leq K}\rho_{i}\rho_{j}(\frac{1}{x_{i}}-\frac{1}{x_{j}})\Bigl(\frac{f(x_{i})}{1-x_{i}}-\frac{f(x_{j})}{1-x_{j}}\Bigr).$
@@
When $K=2$, denote $x=x_{1}$ and $y=x_{2}$ with $0<x<1<y$. Solving Eq. 66, we obtain $\rho_{1}=\frac{x(y-1)}{y-x}$ and $\rho_{2}=\frac{y(1-x)}{y-x}$. Then Eq. 69 becomes to
@@
(70) $\begin{split}\exists\ \alpha>0\textrm{ s.t. }&\bigl(f^{\prime}(y)-f^{\prime}(x)\bigr)^{2}\geq\alpha(\frac{1}{x}-\frac{1}{y})\Bigl(\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\Bigr),\ \forall\ 0<x<1<y\end{split}$
@@
with $\alpha=\alpha_{f}$. This is the same as 4.2. It represents a special case for the discrete inequality Eq. 65 with $K=2$ and thus provides a necessary condition for the discrete inequality. It is also necessary for the continuous inequality Eq. 55 by smoothing the indicator functions. It remains to show that 4.2 is also sufficient, i.e., the case $K=2$ holds implies that the inequality Eq. 69 holds for any $K\geq 3$.
@@
We first show that cases $K\geq 3$ can be reduced to $K=3$. To do so, we use the original formulation (65). We fix $x_{i}>0$, $\alpha_{f}>0$, and $M=\sum_{i=1}^{K}\rho_{i}f^{\prime}(x_{i})$, and consider the following linear programming problem with respect to variables $\{\rho_{i}\}_{i=1}^{K}$
(71a) $\min$ $\sum_{i=1}^{K}\rho_{i}\Bigl(f^{\prime}(x_{i})^{2}-\alpha_{f}\frac{f(x_{i})}{x_{i}}\Bigr)-M^{2}$
(71b) s.t. $\sum_{i=1}^{K}\rho_{i}=1,\ \sum_{i=1}^{K}\frac{\rho_{i}}{x_{i}}=1,\ \sum_{i=1}^{K}\rho_{i}f^{\prime}(x_{i})=M,\ \rho_{i}\geq 0.$
@@
The linear programming problem has a feasible solution and the constraint set is bounded. We claim that there is a minimizer that has at most $3$ nonzero $\rho_{i}$. Otherwise, without loss of generality, we assume for the minimizer $\rho$, it holds $\rho_{j}>0$ for $1\leq j\leq 4$. We can perturb $\rho$ in the direction $\vec{d}=[d_{1},\,d_{2},\,d_{3},\,d_{4}]$ satisfying $\sum_{j=1}^{K}d_{j}=0$, $\sum_{j=1}^{K}\frac{d_{j}}{x_{j}}=0$, and $\sum_{j=1}^{K}d_{j}f^{\prime}(x_{j})=0$. Since there are three linear equations in four variables, there exists a nonzero solution $\vec{d}$. Then, for any $\epsilon$, the perturbed vector $\rho+\epsilon\vec{d}$ still satisfies the constrains Eq. 71b. By treating both the objective function in Eq. 71a and $\rho$ as functions of $\epsilon$, and choosing $\epsilon$ appropriately (either negative or positive), we can decrease the objective function or increase the number of zero entries in $\rho$. This leads to a contradiction so the claim is true. This means that we only need to prove the inequality for the case $K=3$. Using the equivalent formulation (69), assuming (59) holds, we need to show
(72) $\sum_{1\leq i,j\leq 3}\rho_{i}\rho_{j}\Bigl(f^{\prime}(x_{i})-f^{\prime}(x_{j})\Bigr)^{2}\geq\alpha_{f}\sum_{1\leq i,j\leq 3}\rho_{i}\rho_{j}(\frac{1}{x_{i}}-\frac{1}{x_{j}})\Bigl(\frac{f(x_{i})}{1-x_{i}}-\frac{f(x_{j})}{1-x_{j}}\Bigr).$
@@
The last step is to use the inequality in the case of $K=2$ (59) to prove (72). We have the following two technical lemmas whose proofs are in Appendix A.
@@
Given a convex function $f\in C^{2}(0,+\infty)$ with $f(1)=0$ and $f^{\prime}(1)=0$, assume that 4.2 holds. Then there exists $0<\delta<1$, and $0<\alpha_{\delta}\leq\alpha$, such that
(73) $\bigl(f^{\prime}(y)-f^{\prime}(x)\bigr)^{2}\geq\alpha_{\delta}(\frac{1}{x}-\frac{1}{y})\Bigl(\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\Bigr),\ \forall\ x\in(1-\delta,1+\delta)\ \textrm{and}\ y>0.$
@@
Given a convex function $f\in C^{2}(0,+\infty)$ with $f(1)=0$ and $f^{\prime}(1)=0$ and under the constraint (66), conditions (59) and (73) implies (72).
With the above lemmas, the sufficient condition part is proved.
4.3 An easier-to-verify sufficient condition
We also provide a simpler and easier-to-verify sufficient condition on $f$ that guarantees gradient dominance 4.1 holds.
Proposition 4.8.
Assume $f\in C^{2}(0,+\infty)$ is a convex function with $f(1)=0$ and there exists $\alpha_{s}>0$ such that $x^{2}f^{\prime\prime}(x)>\alpha_{s}$ for $0<x\leq 1$. Then 4.2 holds. Note that the assumptions are satisfied with $f^{\prime\prime}(x)=x^{p}$ and $p\leq-2$.
Proof 4.9.
The assumption $x^{2}f^{\prime\prime}(x)>\alpha_{s}$ for $0<x\leq 1$ implies that $f^{\prime\prime}(1)>\alpha_{s}>0$. Combining this with the condition $f\in C^{2}$ ensures the existence of $0<\delta<\frac{1}{2}$, such that
(74) $\frac{f^{\prime\prime}(1)}{2}\leq f^{\prime\prime}(x)\quad\text{for}\quad 1-2\delta\leq x\leq 1+2\delta.$